you kick a soccer ball whose mass is 0.43kgand your foot is in contact with the ball for 8.0ms.The ball leaves your foot with an initial speed of 25m/s.Find the magnitude of impluse and average force exerted by your foot
1
Expert's answer
2018-04-29T13:51:08-0400
Magnitude of impluse:
p=m·v=0.43·25=10.75 (kg·m)/s (N·s)
Average force exerted by your foot
F·Δt=m·v_1-m·v_0
v_0=0
F_average=(mv_1)/Δt=(0.43·25)/(8·〖10〗^(-3) )=1344 N